Circular No. 7784

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 1999gw IN UGC 4881 G. Cresci, Universitá degli Studi, Florence; F. Mannucci, Consiglio

Nazionale delle Ricerche, Florence; R. Maiolino and M. Della Valle, Osservatorio Astrofisico di Arcetri, Florence; and F. Ghinassi, Telescopio Nazionale Galileo (TNG), report the discovery of an apparent supernova during a campaign of near-infrared monitoring of luminous infrared galaxies. SN 1999gw was discovered at mag 17.8 in a K_{s}-band image obtained on 1999 Dec. 16.15 UT with the TNG (+ ARNICA), when compared with an image taken on 2001 Aug. 2.15 with the same telescope (+ NICS). SN 1999gw is located at α = 9^{h}15^{m}54^{s}.7, δ = +44^{o}19′55′′ (equinox 2000.0), which is 3′′.5 north of the K_{s} nucleus A of UGC 4881 (uncertainty on relative position ∼ 0′′.4). An additional K_{s}-band image taken on 2000 Jan. 26.1 with the TNG (+ ARNICA) shows SN 1999gw at K_{s} = 18.4.

SUPERNOVAE 2001iq AND 2001ir R. Chornock and A. V. Filippenko, University of California, Berkeley,

report that inspection of uncalibrated CCD spectra (range 330–1000 nm) obtained on 2001 Dec. 23 UT with the Keck-I telescope reveals that SN 2001iq (IAUC 7780) is of type Ia and near maximum brightness. SN 2001ir (IAUC 7781) is of type IIn; the spectrum is dominated by relatively narrow hydrogen Balmer emission lines.

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001iq, obtained by M. Calkins on 2001 Dec. 23.08 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a reddened type-Ia supernova near maximum light. A recession velocity of 5575 km/s for the host galaxy is derived from narrow emission lines superposed on the spectrum of SN 2001iq; the supernova expansion velocity is ∼ 11 600 km/s for Si II (rest 635.5 nm). The spectral-feature age (cf. IAUC 7774) of SN 2001iq is 1 ± 2 days before maximum light. The spectrum shows strong narrow Na I D absorptions at zero velocity (0.13 nm equivalent width) and at the velocity of the host galaxy (0.36 nm EW). The Schlegel et al. (1998, Ap.J. 500, 525) dust maps indicate a Galactic reddening of E(B − V) = 0.14. Also, the overall shape of the spectrum declines toward the blue, implying that SN 2001iq is heavily extinguished by dust. A spectrum of SN 2001ir, obtained by Calkins on Dec. 24.38 UT, shows it to be a type-IIn supernova, with strong H Balmer emission lines superposed on a nearly featureless blue continuum.

2002 January 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green